A student follows the procedure described in the experiment and collected the following data: 25.5 mL of hydrogen gas collected, barometric pressure of 755 mm Hg, and a temperature of 298 K. Use the ideal gas law (PV=nRT) to answer the questions below. The R value is 0.08206 L atm / Kmol and our data must be in these units. The temperature is given in Kelvin, so no conversion is needed. But the hydrogen gas volume and pressure need conversion and/or correction.  

1) The volume collected is given as 25.5 mL. What is this in liters (L)? Write this number and unit in the #1 space. _______________________

2) The pressure needs to be corrected for the pressure of water vapor. Use Table 1 on page 2 of the lab to look this up and write the water vapor pressure and unit in the #2 space. _________________

3) Subtract the water vapor pressure from the barometric pressure. Write that number and unit in the #3 space. _________________

4) Convert the pressure from #3 from mm Hg to atmospheres and write it in the #4 space. Remember that 1 atm=760 mm Hg. Be sure to write the unit. __________

5) Now that all the needed numbers are corrected and in the proper units, insert them into the ideal gas law equation and solve for moles. Put that answer in space #5. Be sure to write the unit. _____________________

6) Knowing the mass of the unknown metal that was reacted, and using the moles calculated above, the molecular mass of the unknown metal can be calculated. (Recall from the reactions on page 2 that moles of gas = moles of metal). Write the molecular mass in space #6. Be sure to write the unit. __________________

7) Using the molar masses from the periodic chart, identify the unknown metal and write its name in Space #7. If you got the answer Zn, you are correct. ________________
Answer #1

1. volume in liter = 25.5/1000 = 0.0255 L

2. water vapour pressure = 23.8 mmHg

3. remaining pressure = 755 mm Hg- 23.8 mmHg = 731.2 mm Hg

4.   one atm = 760 mm Hg

so 731.2 mm Hg = 0.962 atm

5. mole of ideal gas is = 0.962*0.0255/0.08206*298

=0.001 mole

6. mass of unknown metal that was reacred was not given.

if we assume that mass = 0.065 g

tham molar mass = 0.065/0.001 = 65 g/mol

7. tis molar mass equal to molar mass if Zn (65.38g/mol).